Wednesday, 26 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

Christopher Deninger, Universität Münster, Münster, Germany Some Analogies Between Number Theory and Dynamical Systems on Foliated Spaces (I: p.5)

Chair: Gerd Faltings, Max-Planck-Institut für Mathematik, Bonn, Germany

10:15 – 11:15  Auditorium Maximum H 105

Giovanni Gallavotti, Universitá La Sapienza, Roma, Italy The Chaotic Hypothesis and Universal Large Deviations Properties (I: p.6)

Chair: Rudolf Seiler, Technische Universität, Berlin, Germany

11:45 – 12:45  Auditorium Maximum H 105

Jean-Michel Bismut, Université Paris-Sud, Orsay, France Local Index Theory and Higher Analytic Torsion (I: p.5)

Chair: Jean-Pierre Bourguignon, Inst. des Hautes Etudes Scientifques, Bures-sur-Yvette, France

ICM’98 – Program  1